# GAITEX: Human motion dataset from impaired gait and rehabilitation exercises of inertial and optical sensor data

### v1

@@ Abstract
Wearable inertial measurement units (IMUs) offer a cost-effective and scalable means to assess human movement quality in clinical and everyday settings. However, the development of robust sensor-based classification models for physiotherapeutic exercises and gait analysis requires large, diverse datasets, which are costly and time-consuming to collect. Here, we present a multimodal dataset of physiotherapeutic exercises - including correct and clinically relevant variants - and gait-related exercises - including both normal and impaired gait patterns - recorded from 19 participants using synchronized IMUs and marker-based motion capture (MoCap). The dataset includes raw data from nine IMUs and thirty-five optical markers capturing full-body kinematics. Each IMU is additionally equipped with four optical markers, enabling precise comparison between IMU-derived orientation estimates and reference values from the MoCap system. To support further analysis, we also provide processed IMU orientations aligned with common segment coordinate systems, subject-specific OpenSim models, inverse kinematics results, and tools for visualizing IMU orientations in the musculoskeletal context. Detailed annotations of movement execution quality and time-stamped segmentations support diverse analysis goals. This dataset supports the development and benchmarking of machine learning models for tasks such as automatic exercise evaluation, gait analysis, temporal activity segmentation, and biomechanical parameter estimation. To facilitate reproducibility, we provide code for postprocessing, sensor-to-segment alignment, inverse kinematics computation, and technical validation. This resource is intended to accelerate research in machine learning-driven human movement analysis.
@@
To address this gap, we introduce a dataset comprising two complementary recording scenarios. The first scenario focuses on physiotherapeutic exercises commonly used in the treatment of foot drop, a condition characterized by difficulty in lifting the front part of the foot. Healthy participants performed two commonly used rehabilitation exercises under supervision: (1) resisted dorsiflexion and (2) resisted gait simulation. Both exercises were executed in four distinct ways; one correct and three intentionally incorrect variations, which were executed under the guidance of an instructor. This structure provides a controlled framework for the development and evaluation of classification models that can distinguish between correct and different types of incorrect exercise executions, an essential step in building automated feedback systems. Notably, the dataset does not include clinical populations, its primary aim is to support methodological development rather than clinical assessment. The second scenario comprises treadmill walking data designed to investigate gait anomalies. Participants first walked freely on a treadmill at a self-selected pace that increased incrementally by 0.5 km/h per minute, over a total of three minutes. Afterwards, they were then fitted with a knee orthosis that restricted knee flexion to 0° and repeated the walking trial. The comparison of natural and constrained gait enables the study of altered biomechanical patterns, which may inform models for anomaly detection or aid design in wearable assistive technologies.
@@
A key feature of this dataset is the synchronized recording of nine body worn inertial measurement units (IMUs) (Xsens Awinda, Movella Holdings Inc., Henderson, Nevada, United States of America) alongside a state of the art optical Motion Capturing (MoCap) system (Qualisys, Göteborg, Sweden). The MoCap system was used to track the subjects full-body motion as well as the movement of the IMU housings. This setup enables direct validation of IMU-derived motion data against ground truth kinematic information obtained from the optical system. Several research groups have previously employed similar methodologies for dataset acquisition. For instance, Santos et al.[4] recorded gait information with a MoCap system and 2 IMUs. Similarly, Riglet et al.[5] measured gait on normal ground and a treadmill with varying speed using IMUs, a MoCap system and ground reaction force plates.
In addition to the raw recordings, we provide a post-processed version of the IMU data aligned to subject-specific segment coordinate frames. Using OpenSim [6], we further derived joint kinematics via an inverse kinematics workflow based exclusively on IMU orientations, thereby facilitating biomechanically interpretable representations and enabling quality control through visualization. In contrast to the related work by by Scherpereel et al.[7], who relied on optical markers for kinematic estimation, our approach leverages IMU-based inverse kinematics directly.
@@
  • The release of a multimodal dataset comprising raw orientation data from nine body-worn IMUs; time-synchronized MoCap data that track both the participant’s body segments and the IMU sensor housings and manually curated temporal annotations delineating individual repetitions of each task from the continuous sensor streams.
@@
  • The provision of personalized, anatomically scaled musculoskeletal models and corresponding inverse kinematics results derived from the IMU signals, enabling joint angle estimation and visualization of IMU orientations in the biomechanical context.
@@ Methods
In total, 19 healthy adult volunteers participated in the present study, comprising 15 males and 4 females. The data collection campaign was conducted between September 1st and December 15th, 2024. Participants’ ages ranged from 22 to 55 years (mean: 38.16 $\pm$ 10.78 years). Body weight varied between 52 kg and 102 kg (mean: 74.79 $\pm$ 13.19 kg), and participant height ranged from 158 cm to 192 cm (mean: 177.00 $\pm$ 9.20 cm). The Body Mass Index (BMI) of participants varied between 19.57 and 29.39 kg/m² (mean: 23.72 $\pm$ 2.72 kg/m²).
@@
To achieve comprehensive full-body kinematics, 35 reflective markers were attached to each participant (see Figure 2). The marker arrangement was based on the established Institute of Orthopaedic Research (IOR) marker set, slightly adapted for our specific application needs. These adaptations involved the removal of several markers that conflicted with the placement of IMUs (markers on the toes and markers on the lower back) or essential safety equipment (markers on the upper back the sternum and the fingers), preventing their proper attachment.
The Qualisys MoCap system recorded the spatial trajectories of these markers with the eight mentioned infrared cameras positioned around the participants, operating at a sampling frequency of 100 Hz using the QTM software (v2023.3). Through careful calibration and an appropriate setup, we ensured that the marker positions were recorded with an accuracy of less than one millimeter, ensuring high-quality reference data for subsequent analysis.
Additionally, two RGB cameras recorded the performed exercises at 25 frames per second (fps). The video recordings from these RGB cameras were used later to segment the individual repetitions of exercises and to precisely define the start and end points of individual repetition.
@@
To enable validation of the IMU orientation estimates, a custom sensor mount was designed to attach four reflective Qualisys markers directly to each IMU (see Figure 2). This configuration allowed the IMU orientation to be independently derived from the optical motion capture system, facilitating a comparative analysis of IMU-based and marker-based orientation estimates.
@@
Figure 2: Anterior (left) and posterior (right) views of the experimental marker setup. IMUs are shown in blue, Qualisys markers used for segment pose tracking are marked in red, and Qualisys markers mounted on IMUs for optical orientation validation are shown in orange. Marker and sensor labels match the naming conventions used in the dataset.
@@
Over a period of four months, 19 participants performed two physiotherapeutic and two gait-related movement tasks while equipped with the described sensor setup.
The physiotherapeutic exercises were selected from treatments commonly employed for addressing foot drop. The first physiotherapeutic exercise "Resisted Dorsiflexion" (RD) aimed at strengthening the muscles involved, particularly emphasizing the dorsiflexion of the foot. An adjustable resistance band was used to tailor the training intensity individually (depicted in Figure 3).
@@
To gather sufficient data for identifying typical errors in exercise execution, healthy participants without foot drop performed four variations of each exercise during this campaign. One variation corresponded to the correct (textbook) execution, while the other three represented common movement errors observed in patients with foot drop. The individual variants are shown in Figure 3 and are further described in Table 1. Prior to data collection, participants received structured instruction on all four movement variants. Each variant was first demonstrated by the experimenter, after which participants were given the opportunity to practice the corresponding movement under supervision. During this familiarization phase, participants received immediate corrective feedback whenever deviations from the intended execution occurred. Recordings were only initiated once participants demonstrated consistent and reliable performance of each variant.
Figure 3: Overview of the physiotherapeutic exercises and their execution variants. The first exercise RD is illustrated in the upper row, and the second exercise RGS is illustrated in the lower row. For each exercise, the first image from the left illustrates the standardized start position. The second image depicts the correct execution according to physiotherapeutic guidelines. The remaining three images on the right display common erroneous movement patterns observed in individuals with foot drop. Detailed descriptions of the respective deviations can be found in Table 1.
The first gait-related task "Normal Gait" (NG) involved capturing participants’ natural walking patterns on a treadmill at three different speeds. Given the absence of physical impairments, a natural gait pattern with healthy variations was expected. Each participant performed the walking task at three self-selected treadmill speeds, recorded consecutively for approximately one minute each. Starting from an individually chosen baseline speed, the treadmill velocity was increased twice in small increments, resulting in a total speed difference of approximately 1 km/h between the slowest and fastest condition. The exact duration and magnitude of each increment varied across participants.
@@
Notably, temporal synchronization between the Qualisys MoCap system and the Xsens IMUs is acheived by hardware-synchronization using dedicated synchronization units. This ensures temporal alignment during recording.
@@
The assignment of exercise repetitions to its corresponding variant (e.g. correct, toe lifted, etc.) was based on the verbal instructions provided during acquisition and subsequently verified through manual inspection of the video data.
@@
The dataset presented in this study is openly available via Zenodo under the accession number [LINK EINFÜGEN]. Its directory structure is depicted in Figure 6. For each recorded subject, a pseudonym (e.g., austra, darryl) was assigned and used to name a dedicated folder containing the respective data. Within each subject-specific folder, four subfolders were created, each corresponding to one of the exercises performed during the measurement campaign. These subfolders are named using abbreviations of the respective exercise labels (rd, rgs, ng and gwo). The corresponding exercises are explained in detail in Table 1.
@@
To verify the integrity of the dataset, we conducted a comprehensive technical validation addressing three central objectives: (i) to assess whether the applied transformation correctly maps each IMU’s orientation from its local CS into the OpenSim global reference frame; (ii) to evaluate the temporal synchronization between the Xsens and Qualisys systems; and (iii) to quantify how the orientation estimates provided by the Xsens system deviate over time from the marker-based orientation of each IMU, as determined via the Qualisys system. This analysis thus provides a robust validation of both spatial and temporal alignment and allows users to track potential drift or systematic discrepancies in orientation estimation throughout each recording.
@@
In addition, one specific technical issue should be carefully considered when working with the raw IMU orientation data. Post-hoc analysis revealed a deviation in the magnetometer calibration of two IMUs (XSens_Lower_Leg_Right and XSens_Hand_Right). In contrast to the other units, this IMU used a different reference frame for heading estimation, resulting in a systematic yaw offset. This discrepancy is accounted for in the transformation step ${}^{OpenSim_{y}}_{OpenSim}T$ (as defined in 1) applied during the processing pipeline. Users intending to analyze raw orientations directly are advised to apply the corresponding correction before performing cross-sensor comparisons involving this device.

### v2

# GAITEX: Human motion dataset of impaired gait and rehabilitation exercises using inertial and optical sensors

@@ Abstract
Wearable inertial measurement units (IMUs) provide a cost-effective approach to assessing human movement in clinical and everyday environments. However, developing the associated classification models for robust assessment of physiotherapeutic exercise and gait analysis requires large, diverse datasets that are costly and time-consuming to collect. We present a multimodal dataset of physiotherapeutic and gait-related exercises, including correct and clinically relevant variants, recorded from 19 healthy subjects using synchronized IMUs and optical marker-based motion capture (MoCap). It contains data from nine IMUs and 68 markers tracking full-body kinematics. Four markers per IMU allow direct comparison between IMU- and MoCap-derived orientations. We additionally provide processed IMU orientations aligned to common segment coordinate systems, subject-specific OpenSim models, inverse kinematics outputs, and visualization tools for IMU-derived orientations. The dataset is fully annotated with movement quality ratings and timestamped segmentations. It supports various machine learning tasks such as exercise evaluation, gait classification, temporal segmentation, and biomechanical parameter estimation. Code for postprocessing, alignment, inverse kinematics, and technical validation is provided to promote reproducibility.
@@
To address this gap, we introduce a dataset comprising two complementary recording scenarios. The first scenario focuses on physiotherapeutic exercises commonly used in the treatment of foot drop, a condition characterized by difficulty in lifting the front part of the foot. Healthy participants performed two commonly used rehabilitation exercises under supervision: (1) resisted dorsiflexion and (2) resisted gait simulation. Both exercises were executed in four distinct ways; one correct and three intentionally incorrect variations, which were executed under the guidance of an instructor. This structure provides a controlled framework for the development and evaluation of classification models that can distinguish between correct and different types of incorrect exercise executions, an essential step in building automated feedback systems. Notably, the dataset does not include individuals from clinical populations, as its primary aim is to support methodological development rather than clinical assessment. The second scenario comprises treadmill walking data designed to investigate gait anomalies. Participants first walked freely on a treadmill at a self-selected pace that increased incrementally by 0.5 km/h per minute, over a total of three minutes. Afterwards, they were then fitted with a knee orthosis that restricted knee flexion to 0° and repeated the walking trial. The comparison of natural and constrained gait enables the study of altered biomechanical patterns, which may inform models for anomaly detection or aid design in wearable assistive technologies.
@@
A key feature of this dataset is the synchronized recording of nine body worn inertial measurement units (IMUs) (Xsens Awinda, Movella Holdings Inc., Henderson, Nevada, United States of America) alongside a state of the art optical Motion Capturing (MoCap) system (Qualisys, Göteborg, Sweden). The MoCap system was used to track the participants’ full-body kinematics as well as the movement of the each IMU housing. This setup enables frame-synchronized, segment-level validation of IMU-derived motion data against marker-based reference trajectories. Several research groups have previously employed similar methodologies, but typically under more constrained conditions, with reduced sensor configurations or with a focus on only gait related exercises. For instance, Santos et al.[4] recorded gait information using a MoCap system and only two IMUs, both attached to the same segment, thereby limiting the spatial information to effectively a single measurement point. The study focuses exclusively on healthy gait and therefore offer valuable insights into locomotor behavior. Our dataset extends these efforts by incorporating physiotherapeutic exercises and gait patterns affected by functional impairments.
In addition to the raw recordings, we provide a post-processed version of the IMU data, in which sensor orientations are expressed relative to subject-specific anatomical segment frames. Using OpenSim [5], we further derived joint kinematics via an inverse kinematics workflow based exclusively on these IMU orientations. This workflow not only yields biomechanically interpretable joint trajectories but also facilitates visual quality control of the reconstructed sensor orientations, enabling researchers to assess the plausibility and internal consistency of the IMU-based data. In contrast to Scherpereel et al.[6], who applied a similar inverse kinematics procedure based on optical marker trajectories, our approach relies solely on wearable sensor data and therefore provides a valuable resource for researchers developing and evaluating IMU-only methods.
@@
  • The release of a multimodal dataset [7] comprising raw orientation data from nine body-worn IMUs; time-synchronized MoCap data that track both the participant’s body segments and the IMU sensor housings and manually curated temporal annotations delineating individual repetitions of each task from the continuous sensor streams.
@@
  • The provision of personalized, anatomically scaled musculoskeletal models and corresponding inverse kinematics results derived from the IMU signals, enabling joint angle estimation and visualization of IMU orientations.
@@ Methods
In total, 19 adult volunteers (15 male, 4 female) participated in the present study. All participants were free from any known allergic reactions to Kinesio tapes or adhesives, did not present with acute general illnesses or orthopedic conditions incompatible with participation, and were not pregnant at the time of data acquisition. The data collection campaign was conducted between September 1st and December 15th, 2024. Participants’ ages ranged from 22 to 55 years. For men, the mean age was 39.07 ± 10.44 years (range: 22–55 years), and for women, 34.75 ± 12.97 years (range: 25–53 years). Participant height ranged from 158 cm to 192 cm. Men had a mean height of 180.53 ± 6.33 cm (range: 171–192 cm), while women had a mean height of 163.75 ± 4.92 cm (range: 158–170 cm). Body weight varied between 52 kg and 102 kg. The mean body weight was 78.40 ± 11.33 kg for men (range: 62–102 kg) and 61.25 ± 11.64 kg for women (range: 52–78 kg). The Body Mass Index (BMI) ranged from 19.57 to 29.39 kg/m². For men, the mean BMI was 23.98 ± 2.58 kg/m² (range: 20.91–29.39 kg/m²), and for women, 22.76 ± 3.42 kg/m² (range: 19.57–26.99 kg/m²).
@@
To achieve comprehensive full-body kinematics, 33 reflective markers were attached to each participant (see Figure 2). The marker arrangement was based on the established Institute of Orthopaedic Research (IOR) marker set, slightly adapted for our specific application needs. These adaptations involved the removal of several markers that conflicted with the placement of IMUs (markers on the toes and markers on the lower back) or essential safety equipment (markers on the upper back the sternum and the fingers), preventing their proper attachment.
The Qualisys MoCap system recorded the spatial trajectories of these markers with the eight mentioned infrared cameras positioned around the participants, operating at a sampling frequency of 100 Hz using the QTM software (v2023.3). To ensure high-quality reference data for subsequent analysis, we carefully calibrated the Qualisys motion capture system following the manufacturer’s recommendations. Calibration was only accepted if the standard deviation of wand length measurements remained below 1 mm, indicating stable camera positioning and consistent marker detection across views.
Additionally, two RGB cameras recorded the performed exercises at 25 frames per second. The video recordings from these RGB cameras were used later to segment the individual repetitions of exercises and to precisely define the start and end points of individual repetition.
@@
To enable validation of the IMU orientation estimates, a custom sensor mount was designed to attach four reflective Qualisys markers directly to each IMU. The IMU on the right forefoot carried only three markers due to space constraints (see Figure 2). This configuration allowed the IMU orientation to be independently derived from the optical motion capture system, facilitating a comparative analysis of IMU-based and marker-based orientation estimates.
@@
Figure 2: Anterior (left) and posterior (right) views of the experimental marker setup. IMUs are shown in blue, Qualisys markers used for segment pose tracking are marked in red, and Qualisys markers mounted on IMUs for optical orientation validation are shown in orange. Marker and sensor labels match the naming conventions used in the dataset. Note that the IMU placed on the right forefoot is labeled as XSens_Hand_Right, since the standard Xsens setup does not include a dedicated IMU for this position. Therefore, the otherwise unused right-hand IMU was repurposed for the right forefoot.
@@
Over a period of four months, 19 participants performed two physiotherapeutic and two gait-related movement tasks while equipped with the described sensor setup. Each participant performed the movement tasks in a predefined, non-randomised order, which corresponds to the sequence listed in Table 1.
The physiotherapeutic exercises were selected from treatments commonly employed for addressing foot drop. The first physiotherapeutic exercise "Resisted Dorsiflexion" (RD) aimed at strengthening the muscles involved, particularly emphasizing the dorsiflexion of the foot. An adjustable resistance band was used to tailor the training intensity for each participant (depicted in Figure 3). The tension was adjusted according to the participant’s individual strength and range of motion by the supervisors.
@@
To gather sufficient data for identifying typical errors in exercise execution, healthy participants without foot drop performed four variations of each exercise during this campaign. One variation corresponded to the correct (textbook) execution, while the other three represented common movement errors observed in patients with foot drop. The individual variants are shown in Figure 3 and are further described in Table 1. Prior to data collection, participants received structured instruction on all four movement variants. Each variant was first demonstrated by the experimenter, after which participants were given the opportunity to practice the corresponding movement under supervision. During this familiarization phase, participants received immediate corrective feedback whenever deviations from the intended execution occurred. Recordings were only initiated once participants demonstrated consistent and reliable performance of each variant. To minimise the potential influence of fatigue, participants were instructed to take a brief rest after completing each exercise variant, corresponding to ten repetitions per set.
Figure 3: Overview of the physiotherapeutic exercises and their execution variants. Both exercises are performed in a standing position. The first exercise RD is illustrated in the upper row, and the second exercise RGS is illustrated in the lower row. For each exercise, the first image from the left illustrates the standardized start position. The second image depicts the correct execution according to physiotherapeutic guidelines. The remaining three images on the right display common erroneous movement patterns observed in individuals with foot drop. Detailed descriptions of the respective deviations can be found in Table 1.
The first gait-related task "Normal Gait" (NG) involved capturing participants’ natural walking patterns on a treadmill at three different speeds. Given the absence of physical impairments, a natural gait pattern with healthy variations was expected. Each participant performed the walking task at three self-selected treadmill speeds, recorded consecutively for approximately one minute each. Starting from an individually chosen baseline speed, the treadmill velocity was increased twice in small increments, resulting in a total speed difference of approximately 1 km/h between the slowest and fastest condition. The exact duration and magnitude of each increment varied across participants and can be retrieved from the corresponding timestamps_... file (see Figure 6).
@@
Notably, temporal synchronization between the Qualisys MoCap system and the Xsens IMUs is achieved by hardware-synchronization using dedicated synchronization units. This ensures temporal alignment during recording.
@@
The assignment of exercise repetitions to its corresponding variant (e.g. correct, toe lifted, etc.) was based on the verbal instructions given to the participants during data acquisition and subsequently verified through manual inspection of the video data by a supervisor. Since the labels were derived directly from the instructed exercise variants rather than from independent manual annotation, the labels should be interpreted as reflecting the intended exercise condition rather than a qualitative assessment of the movement execution.
@@
The dataset [7] presented in this study is openly available via Zenodo. Its directory structure is depicted in Figure 6. For each recorded subject, a pseudonym (e.g., austra, darryl) was assigned and used to name a dedicated folder containing the respective data. Within each subject-specific folder, four subfolders were created, each corresponding to one of the exercises performed during the measurement campaign. These subfolders are named using abbreviations of the respective exercise labels (rd, rgs, ng and gwo). The corresponding exercises are explained in detail in Table 1.
@@
To verify the integrity of the dataset, we conducted a technical validation addressing three central objectives: (i) to assess whether the applied transformation correctly maps each IMU’s orientation from its local CS into the OpenSim global reference frame; (ii) to offer exemplary support for a successful temporal synchronization between the Xsens and Qualisys systems and (iii) to quantify how the orientation estimates provided by the Xsens system deviate over time from the marker-based orientation of each IMU, as determined via the Qualisys system. This analysis thus provides a robust validation of both spatial and temporal alignment and allows users to track potential drift or systematic discrepancies in orientation estimation throughout each recording.
@@
In addition, one specific technical issue should be carefully considered when working with the IMU orientation data. Post-hoc analysis revealed a deviation in the magnetometer calibration of two IMUs (XSens_Lower_Leg_Right and XSens_Hand_Right). In contrast to the other units, this IMU used a different reference frame for heading estimation, resulting in a systematic yaw offset. This discrepancy is accounted for in the transformation step $^{OpenSim_{y}}_{OpenSim}T$ (as defined in equation 1) applied during the processing pipeline. Accordingly, this issue is already compensated for in the processed orientation trajectories (xsens_imu_data_segment_registered_<subject>_<exercise_abbreviation>.csv), but not in the raw data (xsens_imu_data_<subject>_<exercise_abbreviation>.csv). The required compensation transformation $^{OpenSim_{y}}_{OpenSim}T$ is computed individually for each recording. The exact value of this deviation can be determined for every measurement using the provided raw data and the processing scripts available in the associated repository.